18/135 and 90/675 are proper, unlike and                     fractions.
  18/135 and 90/675 are proper, unlike and                     fractions.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

18/135 and 90/675 are proper, unlike and equivalent fractions. Consider the two given fraction, 18/135 and 90/675

So,
(18/135) = (90/675)
By cross multiplication,
we get
(18 × 675) = (90 × 135)
12,150 = 12,150
Therefore, 18/135 and 90/675 are proper, unlike and equivalent fraction




Thanks